### v1

# CGCE [Uncaptioned image]: Classifier-Guided Concept Erasure in Generative Models

Abstract
Recent advancements in large-scale generative models have enabled the creation of high-quality images and videos, but have also raised significant safety concerns regarding the generation of unsafe content. To mitigate this, concept erasure methods have been developed to remove undesirable concepts from pre-trained models. However, existing methods remain vulnerable to adversarial attacks that can regenerate the erased content. Moreover, achieving robust erasure often degrades the model’s generative quality for safe, unrelated concepts, creating a difficult trade-off between safety and performance. To address this challenge, we introduce Classifier-Guided Concept Erasure (CGCE), an efficient plug-and-play framework that provides robust concept erasure for diverse generative models without altering their original weights. CGCE uses a lightweight classifier operating on text embeddings to first detect and then refine prompts containing undesired concepts. This approach is highly scalable, allowing for multi-concept erasure by aggregating guidance from several classifiers. By modifying only unsafe embeddings at inference time, our method prevents harmful content generation while preserving the model’s original quality on benign prompts. Extensive experiments show that CGCE achieves state-of-the-art robustness against a wide range of red-teaming attacks. Our approach also maintains high generative utility, demonstrating a superior balance between safety and performance. We showcase the versatility of CGCE through its successful application to various modern T2I and T2V models, establishing it as a practical and effective solution for safe generative AI.
Figure 1: We present CGCE, an efficient plug-and-play framework for robust and high-fidelity concept erasure. Top: CGCE produces safer and higher-quality results compared to state-of-the-art baselines [8, 10, 38, 30] across diverse T2I erasure tasks, including nudity, artistic style, and object removal. Bottom: The cross-modal safety and versatility of CGCE, which can be seamlessly applied as a safeguard to a range of modern T2I and T2V models to ensure safe generation without altering their original weights. Sensitive content (*) has been masked for publication.
1 Introduction
Recent advancements in large-scale diffusion models have led to remarkable success in Text-to-Image (T2I) [27, 23, 7, 18, 11] and Text-to-Video (T2V) [36, 17, 15] generation, producing high-fidelity and diverse visual content from simple text descriptions. These models have demonstrated an impressive ability to synthesize a wide range of concepts, unlocking numerous creative and commercial applications. However, the reliance on web-scraped training data [29] poses significant safety and ethical risks, as these datasets often contain harmful, biased, or copyrighted material. A primary concern is the generation of NSFW content, such as sexually explicit imagery, which raises urgent questions about the responsible deployment of these powerful models.
Several strategies have been proposed to mitigate the generation of such undesirable content. A straightforward approach is dataset filtering [1], which involves removing unsafe images before training. However, this is often computationally expensive and impractical for every new concept because aggressive filtering can inadvertently degrade overall generative quality. Post-hoc solutions have emerged as more practical alternatives. Model fine-tuning approaches [8, 12, 40, 16, 30] permanently alter the model’s weights to erase or remap harmful concepts, but this process can be resource-intensive and often leads to a trade-off between erasure effectiveness and utility preservation of safe content. A distinct category is model editing [9, 10] which uses closed-form solutions to directly alter model weights without iterative training. While highly efficient, these edits can also be incomplete and may remain vulnerable to adversarial attacks that regenerate the erased concepts. Finally, training-free, filtering-based methods [28, 38] modify the generation process at inference time by steering the output away from undesired concepts. While flexible, these methods can be less robust and are often limited to specific model architectures or susceptible to being bypassed, especially when prompts express harmful concepts implicitly.
To address these limitations, we introduce Classifier-Guided Concept Erasure (CGCE), an efficient framework that offers a robust and versatile solution for safe visual generation, as shown in Fig. 1. A key strength of CGCE is its plug-and-play nature, allowing it to be applied to diverse generative models, such as modern T2I and T2V architectures, without altering the original model weights. This preserves the high-quality generative capabilities of the base model for safe content, a critical advantage over many fine-tuning methods that suffer from utility degradation. Our approach leverages a lightweight, specially trained classifier that operates in the text embedding space. This classifier is first trained on a synthetic dataset of paired safe and unsafe prompts, enabling it to accurately detect the presence of an undesired concept within a given text embedding. At inference time, our classifier serves a dual role. It first acts as a safeguard, inspecting the text embedding to detect the presence of an undesired concept. Upon detection, it then functions as a refiner, altering the embedding to neutralize the harmful semantics and guide it toward a safe semantic region. This process ensures a safe visual output from the T2I or T2V model while leaving safe prompts untouched, thereby preserving the model’s original generation quality. Furthermore, our framework is highly scalable and can simultaneously erase multiple distinct concepts by aggregating the guidance from several classifiers.
Our extensive experiments demonstrate the effectiveness of CGCE across multiple tasks, including nudity, artistic style, and object removal. CGCE achieves state-of-the-art performance on multiple benchmarks while maintaining the utility of the base model. It provides a superior trade-off between safety and quality compared to existing fine-tuning methods. Furthermore, we demonstrate the versatility of CGCE by applying it to various modern T2I and T2V models. Our main contributions can be summarized as:
  • We propose CGCE, a plug-and-play, image-free framework that leverages a classifier to detect and refine text embeddings for robust concept erasure.
  • We introduce a scalable method for multi-concept erasure, allowing CGCE to efficiently remove multiple unwanted concepts in a single refinement process.
  • Our method achieves state-of-the-art safety performance on multiple benchmarks without degrading the generative quality of the base model, offering a superior trade-off between utility and robustness.
  • We demonstrate CGCE’s versatility by successfully applying it to various modern T2I and T2V models, showcasing its potential as a universal safeguard for generative AI.
Figure 2: Overview of CGCE. Stage 1: LLM is used to create a dataset of paired prompts, each containing a safe prompt and a semantically similar unsafe version. Stage 2: A lightweight classifier is trained on the embeddings of these prompts to distinguish between safe and unsafe content. Stage 3: At inference time, the trained classifier acts as a plug-and-play safeguard. If an input prompt is safe, its embedding is passed directly to the generative model. If unsafe, the classifier then acts as a refiner, using its own gradients to iteratively modify the embedding. This process steers the embedding away from the harmful concept before it is passed to the T2I or T2V model to ensure a safe final output.
2 Related Works
Concept Erasure in T2I. Recent work in concept erasure aims to mitigate risks in T2I models by removing undesired concepts. One major category of methods involves model fine-tuning approaches [8, 12, 40] that map an erasing concept to its desired target. However, as such prompt-based methods are vulnerable to adversarial attacks [2, 39, 42], more robust techniques have been developed. These include incorporating an efficient single-timestep attack into the training loop [16], employing a two-stage framework to first identify vulnerabilities [30], or reformulating unlearning as an image-based preference optimization problem [22]. For greater efficiency, closed-form editing methods [9, 10] directly alter weights without iterative optimization. In contrast, methods [28, 38] with unchanged model weights offer flexibility by intervening at inference time. These interventions include modifying the classifier-free guidance [14] process [28] or identifying a “toxic subspace” and projecting input tokens away from it [38].
Concept Erasure in T2V. While concept erasure is well-explored for images, its application to T2V generation is an emerging field with unique challenges, such as maintaining temporal consistency and adapting to modern architectures that lack cross-attention layers. Training-free methods have been extended to this domain, for example by applying a token-filtering mechanism to videos [38]. More recently, a dedicated training-free T2V framework was proposed that first adjusts prompt embeddings and then guides the latent noise to steer generation away from the target concept [34]. In contrast, unlearning-based approaches [37] have adapted T2I fine-tuning techniques for the video domain, incorporating regularization strategies to localize the erasure and preserve the model’s knowledge of related concepts.
Attacks on Concept Erasure. The effectiveness of erasure methods is evaluated using red-teaming attacks [2, 42, 39, 35] that manipulate model inputs to bypass safety mechanisms. Many of these attacks focus on the text prompt. White-box methods leverage access to the model’s internal states, such as gradients or classification capabilities, to automatically find problematic prompts [2, 42]. Black-box methods extract a representation of the inappropriate concept and then use it to automatically generate attack prompts [39]. More advanced frameworks employ a multimodal attack strategy, for instance by using gradient-based optimization in the textual modality to craft adversarial prompts that maintain semantic intent but evade prompt filters [35].
3 Proposed Method
We propose CGCE, a lightweight, plug-and-play framework for safeguarding T2I and T2V models without altering their original weights. First, we use a Large Language Model (LLM) to generate prompt pairs with and without a target concept to train a lightweight classifier. During inference, this classifier acts as a safeguard by detecting the undesired concept in the text embedding. Upon detection, it refines the embedding by steering it away from the concept’s semantic representation using gradient descent before it is passed to the generative model. This process is illustrated in Fig. 2.
3.1 Concept Classifier
Synthetic Data Generation. Our goal is to build a classifier capable of detecting target concepts within the text embedding space used by T2I and T2V models. These models typically process an input prompt $p$, through a text encoder like CLIP [24] or T5 [25] to generate a text embedding. To train our classifier, we first construct a dataset of prompt pairs using a LLM (e.g., Gemini [3], Qwen [31]). Each pair consists of a prompt containing the target concept (unsafe, denoted $p_{u}$) and a corresponding prompt that preserves similar content but excludes the concept (safe, denoted $p_{s}$). This paired data structure is crucial for isolating the specific semantic features of the target concept, which helps the classifier learn a more precise decision boundary and resolves ambiguity about which features to identify. To enhance the classifier’s robustness, we ensure the prompt dataset is diverse across various scenarios. For instance, when targeting the “nudity” concept, a pair would include an unsafe prompt, $p_{u}$: a photo of a nude girl” and its corresponding safe prompt, $p_{s}$: a photo of a girl”. The templates used for prompting the LLM are detailed in the Appendix for reproducibility.
Design the Classifier. Given an input prompt $p$, a text encoder $\boldsymbol{\mathcal{T}}$ (e.g., CLIP or T5) processes it into a text embedding $\boldsymbol{\varepsilon}_{p}=\boldsymbol{\mathcal{T}}(p)\in\mathbb{R}^{n\times d}$ where $n$ is the number of tokens and $d$ is the embedding dimension. To guide the classifier $\boldsymbol{f}_{\theta}$, we also define a target concept prompt $c$, which is encoded into a concept embedding $\boldsymbol{\varepsilon}_{c}=\boldsymbol{\mathcal{T}}(c)$, where $\boldsymbol{\varepsilon}_{c}\in\mathbb{R}^{m\times d}$ for $m$ concept tokens. For example, to detect nudity, $c$ could be “sexual, nudity, porn, naked”.
Firstly, the text embedding $\boldsymbol{\varepsilon}_{p}$ and $\boldsymbol{\varepsilon}_{c}$ are projected into a lower-dimensional space, followed by a multi-head cross-attention layer [32]. The intuition is to identify which tokens in the input prompt embedding $\boldsymbol{\varepsilon}_{p}$ are most semantically related to the tokens in the concept embedding $\boldsymbol{\varepsilon}_{c}$. The input prompt tokens act as queries that attend to the concept tokens, which serve as both keys and values. This allows the model to focus on the parts of the prompt that are semantically aligned with the concept we want to detect.
Mathematically, given the query matrix $\mathbf{Q}=\boldsymbol{\varepsilon}_{p}\mathbf{W}_{q}$, the key matrix $\mathbf{K}=\boldsymbol{\varepsilon}_{c}\mathbf{W}_{k}$ and the value matrix $\mathbf{V}=\boldsymbol{\varepsilon}_{c}\mathbf{W}_{v}$, the attention weight matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ is computed using the scaled dot-product attention formula:
$\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right),$ (1)
where $d_{k}$ is the dimension of the keys. The output of this layer $\boldsymbol{\varepsilon}_{att}=\mathbf{A}\mathbf{V}$ is a new representation of the prompt embedding, weighted by their relevance to the concept.
To create a single representation for the entire prompt, we first derive an importance score $\boldsymbol{s}\in\mathbb{R}^{n}$, by taking the maximum attention weight that each prompt token $i$ assigned to any of the concept tokens $j$ ($\boldsymbol{s}_{i}=\max_{j}\mathbf{A}_{ij}$). These scores are normalized via a softmax function to produce a final importance distribution $\boldsymbol{\alpha}=\text{softmax}(\boldsymbol{s})$. This distribution is used to compute an aggregated sequence representation $\boldsymbol{\varepsilon}_{seq}$, by taking a weighted sum of the attended features: $\boldsymbol{\varepsilon}_{seq}=\sum_{i=1}^{n}\boldsymbol{\alpha}_{i}\cdot\boldsymbol{\varepsilon}_{att,i}$. This final vector $\boldsymbol{\varepsilon}_{seq}$, which emphasizes the prompt’s concept-related features, is then passed through an MLP and a sigmoid function to yield the final probability $\boldsymbol{f}_{\theta}(\boldsymbol{\varepsilon}_{p},\boldsymbol{\varepsilon}_{c})$, indicating the likelihood that the prompt contains the target concept.
Training Loss. The entire classifier $\boldsymbol{f}_{\theta}$, including the linear projection layers, cross-attention module, and final linear layers, is trained end-to-end using a Binary Cross-Entropy loss. Given our paired dataset of prompts, we assign a ground-truth label $y=1$ for unsafe prompts $p_{u}$ and $y=0$ for the corresponding safe prompts $p_{s}$. The loss function $\mathcal{L}$ is then formulated as the expectation over all prompt embeddings $\boldsymbol{\varepsilon}_{p}$ in the training data:
[…]
where $\hat{y}=\boldsymbol{f}_{\theta}(\boldsymbol{\varepsilon}_{p},\boldsymbol{\varepsilon}_{c})$ is the predicted probability from the classifier.
3.2 Concept Erasure with Classifier
Table 1: Classifier performance using different text encoders. We report the detection accuracy (%) on our unsafe nudity dataset and the false positive rate (%) on the safe COCO-30K dataset.
[…]
CLIP [24] 99.22 7.18
T5-XXL [25] 99.71 11.32
Classifier as a Safeguard. Once trained, the classifier is integrated as a plug-and-play module to safeguard T2I or T2V models at inference time. Given a user prompt $p$ and a target concept prompt $c$, the classifier takes their respective embeddings $\boldsymbol{\varepsilon}_{p}$ and $\boldsymbol{\varepsilon}_{c}$ as input. A concept is detected if the classifier’s output probability $\boldsymbol{f}_{\theta}(\boldsymbol{\varepsilon}_{p},\boldsymbol{\varepsilon}_{c})$, exceeds a predefined threshold $\tau$. In practice, we set the threshold $\tau=0.5$. As shown in Tab. 1, our classifier detects nearly 100% of unsafe content on a large dataset of over 3000 nudity prompts, while maintaining a low false positive rate of just 7.18% and 11.32 % on the safe COCO-30K dataset [19]. If the concept is not detected, the original text embedding $\boldsymbol{\varepsilon}_{p}$ is passed to the generative model without modification. This is a crucial advantage, as it fully preserves the original model’s utility for safe prompts, a property that is often compromised by prior methods that permanently alter model weights.
Classifier as a Refiner. If the target concept is detected in an unsafe prompt $p_{u}$, the classifier then acts as a refiner to steer the embedding $\boldsymbol{\varepsilon}_{p_{u}}$ towards a region of the latent space that does not contain the concept. Intuitively, the objective is to find a refined embedding $\boldsymbol{\varepsilon}^{\prime}_{p_{u}}$ that minimizes the classifier’s prediction, effectively driving its output towards zero. We achieve this by iteratively updating the embedding using weighted gradient descent of the classifier’s output with respect to the embedding itself. The update rule at each refinement step $k$ is given by:
[…]
where $\boldsymbol{\varepsilon}^{(0)}_{p_{u}}=\boldsymbol{\varepsilon}_{p_{u}}$ is the initial unsafe embedding and $\eta$ is the step size. The refinement process is designed to be both precise and stable. First, to make the update more precise, the raw gradient $\nabla\boldsymbol{f}_{\theta}$ is weighted by the token importance score $\boldsymbol{s}^{(k)}$ via an element-wise product $\odot$. This localizes the changes, applying the strongest updates to the tokens most semantically related to the target concept while minimally affecting unrelated parts of the prompt. Second, this weighted gradient $\boldsymbol{g}^{(k)}$ is scaled using a norm-based factor. This factor normalizes the update to prevent excessively large or small steps and then rescales it to be proportional to the magnitude of the embedding, making the overall update stable and adaptive. The embedding is iteratively refined until a stopping condition is met, which is either a fixed number of iterations or the classifier’s prediction falling below a predefined threshold $\tau$. The final refined embedding $\boldsymbol{\varepsilon}^{\prime}_{p_{u}}$ is then passed to the generative model to produce a safe output. This procedure is detailed in the Appendix.
3.3 Multi-Concept Erasure
A significant advantage of our framework is its natural extensibility to simultaneous multi-concept erasure. While many prior methods must be re-applied for each concept, CGCE can be generalized to remove multiple, distinct concepts in a single refinement process.
To erase $L$ distinct concepts, we first train $L$ independent lightweight classifiers $\{\boldsymbol{f}_{\theta_{1}},\boldsymbol{f}_{\theta_{2}},\dots,\boldsymbol{f}_{\theta_{L}}\}$. Each classifier $\boldsymbol{f}_{\theta_{l}}$ is trained using the method from Sec. 3.1 to detect its corresponding target concept $c_{l}$ and its embedding $\boldsymbol{\varepsilon}_{c_{l}}$. During the refinement phase, we generalize the single-concept update rule in Eq. 3. At each iteration $k$, we first compute the weighted gradient $\boldsymbol{g}^{(k)}_{l}$ for each classifier $\boldsymbol{f}_{\theta_{l}}$. These individual weighted gradients are then normalized and summed to create a single aggregated update vector $\boldsymbol{G}^{(k)}$. This aggregated vector is then scaled and applied to update the embedding:
[…]
This aggregation allows CGCE to efficiently steer the embedding away from multiple unwanted semantic regions simultaneously, providing a highly scalable solution for multi-concept erasure.
3.4 Plug-and-Play Safeguarding for Diverse Generative Models
Many existing concept erasure methods are architecture-specific, often developed and tested exclusively on models like Stable Diffusion v1.4 (SD-v1.4) [27]. While training-free approaches like SAFREE [38] offer greater flexibility by operating across diverse model backbones, their reliance on token-level proximity to a predefined “toxic subspace” can limit their effectiveness against prompts where harmful concepts are expressed implicitly or contextually. Our method is designed to address this gap. By training a classifier on a dataset of entire prompt pairs, our approach learns the holistic semantic meaning and context of the prompt, rather than just the proximity of individual words. This enables the classifier to identify unsafe patterns and combinations of words that may only be harmful when used together, providing greater robustness against implicit prompts. Furthermore, since our lightweight classifier operates on the text embedding space, it functions as a plug-and-play module that can be easily integrated with various T2I and T2V models without requiring any modification to their weights. For example, a classifier trained on CLIP embeddings can safeguard both SD-v1.4 and SD-v2.1 [27], while one trained on T5 embeddings is compatible with modern models like SD-v3 [7] and FLUX [18]. To demonstrate our method’s versatility and effectiveness, Sec. 4.3 provides a detailed qualitative and quantitative evaluation on a range of modern generative models, including T2I architectures like SD-v3 [7], FLUX [18] and VAR [33, 11], and T2V models such as HunyuanVideo [17] and CogVideoX [36].
4 Experiments
Table 2: Attack Success Rate (ASR) and generation quality comparison of concept erasure methods for the nudity removal task. All methods are benchmarked on the SD-v1.4 backbone. Bold: best. Underline: second-best.
[…]
No Weights
Modification
I2P $\downarrow$ SixCD $\downarrow$ P4D $\downarrow$ RAB $\downarrow$ MMAD $\downarrow$ UDA $\downarrow$ FID $\downarrow$ CLIP $\uparrow$
SD-v1.4 [27] - 36.41 82.13 75.74 97.19 63.60 97.18 16.79 31.31
ESD [8] $\times$ 8.59 16.70 30.15 56.84 9.20 73.94 15.98 30.42
MACE [20] $\times$ 6.02 8.17 8.09 5.61 4.10 66.20 15.14 29.33
UCE [9] $\times$ 6.23 10.27 18.01 14.39 7.90 73.24 17.74 30.30
RECE [10] $\times$ 6.77 20.66 28.68 12.98 36.10 72.53 16.24 30.95
DUO [22] $\times$ 12.03 24.24 23.90 31.23 28.60 82.39 15.83 31.28
STEREO [30] $\times$ 0.75 4.74 5.15 4.21 7.90 28.17 17.98 30.16
SLD-Medium [28] 26.64 73.16 70.96 97.54 52.30 52.11 20.38 30.71
SLD-Strong [28] 20.84 59.39 58.09 91.58 39.60 34.75 23.19 29.97
SLD-Max [28] 15.68 46.46 40.81 64.21 35.60 21.83 26.63 29.20
SAFREE [38] 12.35 39.18 48.90 74.74 44.70 80.99 19.55 30.69
[…]
4.1 Experimental Setup
Baselines. Following recent works [8, 9, 10, 20, 38, 30], we use SD-v1.4 [27] as our primary T2I model for comparison. We evaluate CGCE against eight baselines, including training-based methods (e.g., ESD [8], DUO [22], STEREO [30]), closed-form editing methods (e.g., UCE [9], MACE [20], RECE [10]), and training-free inference methods (e.g., SLD [28], SAFREE [38]). The evaluation spans three main concept removal tasks:
  • Nudity Removal. All methods are tested on standard user prompts from the I2P [28] and SixCD [26] datasets, as well as a wide range of adversarial attacks, including P4D [2], Ring-A-Bell (RAB) [39], MMA-Diffusion (MMAD) [35] and UnlearnDiffAtk (UDA) [42].
  • Artist Style Removal. Following prior works [8, 10, 38, 30], we select “Van Gogh” as the artistic styles to erase. Additionally, we evaluate each method’s robustness against the UDA [42] attack.
  • Object Removal. For object removal, we test the ability to erase the “church” concept, following the experimental setup of the UDA [42] attack.
Generalization to Other Models. To demonstrate our method’s versatility, we follow the protocol [38] and extend our nudity removal tests to other modern T2I backbones. These include DiT-based models (e.g., SD-v3 [7], FLUX.1-dev [18]) and VAR models (e.g., Infinity-2B [11], Switti-AR [33]). For T2V models, we follow T2VU [37] to test nudity removal on CogVideoX-2B, CogVideoX-5B [36] and HunyuanVideo [17], using the Gen [37] and SafeSora-Sexual [5] benchmarks. For brevity, we refer to these models as CogX-2B, CogX-5B and Hunyuan, respectively.
Evaluation Metrics. We measure two main aspects: the effectiveness of concept removal and the preservation of model utility after erasure.
  • Erasure Effectiveness. For all tasks, we report the Attack Success Rate (ASR), where a lower ASR signifies more effective removal. For nudity removal, we use the NudeNet [21] detector with a threshold of 0.45 to identify inappropriate content. For T2V models, we compute a frame-level ASR, calculated as the total number of unsafe frames divided by the total number of generated frames. For artistic style removal, we use the style classifier from [42] to compute the ASR. We also report the LPIPS score [41] of generated images relative to the original SD output. For object removal, we follow established protocols [8, 30] and use a ResNet-50 ImageNet classifier [6] to compute the ASR.
  • Utility Preservation. To measure how well the models retain their ability to generate normal images, we compute the FID [13] and CLIP Score on the COCO dataset [19] using 30K generated images for all experiments. A lower FID score indicates higher image quality, while a higher CLIP score suggests better text-image alignment.
Implementation Details. For each concept erasing task, we train our classifier on a dataset of 1000 generated prompt pairs. The target concept prompt $c$ is set to “sexual, nudity, sex, porn, naked” for the nudity removal task. For other tasks, such as artistic style or object removal, $c$ is simply the name of the concept (e.g., “Van Gogh” or “church”). The classifier was trained for 10 epochs using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a batch size of 32. The step size $\eta$ for the iterative embedding refinement is a task-specific hyperparameter. For our experiments with the SD-v1.4 model, we empirically set $\eta$ to 1.0, 0.15, and 0.5 for nudity, artistic style, and object removal, respectively. Further implementation details are in the Appendix.
4.2 Main Results
Figure 3: Qualitative evaluation of CGCE’s effectiveness in erasing target concepts while preserving unrelated concepts, compared to baseline methods with SD-v1.4 backbone. Sensitive content (*) has been masked for publication.
Nudity Removal. As shown in Tab. 2 and Fig. 3, CGCE achieves state-of-the-art or highly competitive performance across all six red-teaming benchmarks for nudity removal, without requiring any modification to the model weights. Our approach ranks first on four of the six benchmarks (e.g., SixCD, P4D, RAB, and MMAD). On the I2P and UDA benchmarks, CGCE achieves the second-best ASR. Notably, CGCE offers a significantly better balance between robustness and utility preservation than the top-performing methods on these benchmarks. For example, while STEREO achieves a lower ASR on I2P, it shows a noticeable drop in utility on the COCO dataset compared to CGCE. Similarly, SLD-Max has a lower ASR on UDA but suffers from severe utility degradation. In contrast, CGCE maintains FID and CLIP scores that are almost identical to those of the original SD-v1.4 model, indicating that we achieve strong robustness with almost no loss in the model’s general generative ability. This is because our classifier can accurately identify safe and unsafe content, making no changes to safe samples. Furthermore, CGCE consistently and significantly outperforms SAFREE, another training-free approach that operates on text embeddings, across all benchmarks. This is because SAFREE’s reliance on token-level proximity to a predefined ”toxic subspace” is less effective against prompts where harmful concepts are expressed implicitly or contextually.
Artistic Style and Object Removal. CGCE also shows strong results on artistic style and object removal tasks. Artistic Style Removal: Following [10, 38], we first measure the perceptual distance compared to the original SD model using LPIPS [41]. We report $\text{LPIPS}_{e}$ for erased artist styles and $\text{LPIPS}_{u}$ for unerased styles. The overall effectiveness is measured by their difference, $\text{LPIPS}_{d}=\text{LPIPS}_{e}-\text{LPIPS}_{u}$. As shown in Tab. 3, CGCE achieves the best performance with $\text{LPIPS}_{d}$ values of 0.43 and ASR of 0 for “Van Gogh” style. Crucially, CGCE achieves a perfect $\text{LPIPS}_{u}$ score of 0, showing no negative impact on unrelated artist styles. To further test the robustness of CGCE on the artistic style removal task, we evaluate it against the UDA attack (Tab. 4). CGCE achieves the lowest ASR of 24.00, clearly outperforming all other concept-erasing baselines. Object Removal: As shown in Tab. 5, CGCE again achieves the best ASR of 16.00 against the UDA attack for the “church” concept. In both tasks, this state-of-the-art erasure effectiveness is achieved with negligible loss of utility. The FID and CLIP scores remain consistent with those of the original SD model, showing that our method can precisely erase target concepts without harming the model’s ability to generate other content. Figure 3 provides qualitative comparisons of CGCE with the other baselines.
4.3 Generalization and Extensibility of CGCE
Generalization Across Diverse T2I Architectures. To demonstrate the versatility of CGCE, we extend our evaluation to a range of modern T2I backbones, including DiT-based models (e.g., SD-v3, FLUX.1-dev) and VAR models (e.g., Switti-AR, Infinity-2B). We benchmark against SAFREE, as it is a training-free method that can be directly applied to these architectures, providing a fair comparison. As shown in Fig. 4 and Tab. 6, CGCE demonstrates strong generalization across a variety of modern T2I architectures, consistently outperforming both the base models and SAFREE across all adversarial attacks. This advantage is particularly evident on the FLUX model, a guidance-distilled architecture where SAFREE is less effective when its mechanism does not utilize negative prompting. In contrast, CGCE remains highly effective on this model, reducing the ASR of the RAB attack to nearly zero. These results confirm that CGCE operates effectively as a model-agnostic, plug-and-play safeguard suitable for a diverse range of modern generative models.
Table 3: ASR and LPIPS scores comparison of concept erasure methods for Van Gogh art style removal task.
[…]
ESD [8] 0.39 0.23 0.16 5.00
UCE [9] 0.26 0.16 0.10 20.00
RECE [10] 0.41 0.36 0.05 20.00
SAFREE [38] 0.42 0.38 0.04 5.00
STEREO [38] 0.41 0.27 0.14 10.00
[…]
Table 4: ASR comparison of concept erasure methods for Van Gogh art style removal task under the UDA attack [42].
[…]
Metrics
[…]
SD
[…]
ESD [8]
(ICCV’23)
[…]
UCE [9]
(WACV’24)
[…]
RECE [10]
(ECCV’24)
[…]
SAFREE [38]
(ICLR’25)
[…]
STEREO [30]
(CVPR’25)
[…]
CGCE
[…]
Table 5: ASR comparison of concept erasure methods for Church removal task under the UDA attack [42].
[…]
Metrics
[…]
SD
[…]
ESD [8]
(ICCV’23)
[…]
RECE [10]
(ECCV’24)
[…]
SAFREE [38]
(ICLR’25)
[…]
STEREO [30]
(CVPR’25)
[…]
CGCE
[…]
Figure 4: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with modern T2I and T2V architectures. Sensitive content (*) has been masked for publication.
Table 6: ASR comparsion on T2I architectures.
[…]
SD-v3 [7] 31.90 49.97 69.85 83.86 25.20
SD-v3 + SAFREE [38] 16.40 31.43 41.18 58.10 5.90
[…]
FLUX [18] 38.56 66.02 77.94 97.19 55.20
FLUX + SAFREE [38] 37.81 62.04 74.63 87.89 53.90
[…]
Switti-AR [33] 17.72 43.73 51.84 92.98 28.90
Switti-AR + SAFREE [38] 10.09 25.66 26.47 72.63 16.20
[…]
Infinity-2B [11] 35.23 68.10 72.06 95.44 54.80
Infinity-2B + SAFREE [38] 20.41 41.78 31.66 67.37 31.60
[…]
Table 7: ASR comparsion on T2V architectures.
[…]
Hunyuan [17] 59.87 21.27
Hunyuan + SAFREE [38] 29.79 8.66
Hunyuan + T2VU [37] 9.18 5.25
[…]
CogX-2B [36] 28.10 23.00
CogX-2B + SAFREE [38] 16.39 5.17
CogX-2B + T2VU [37] 10.84 4.70
[…]
CogX-5B [36] 36.43 22.70
CogX-5B + SAFREE [38] 23.61 1.23
CogX-5B + T2VU [37] 10.59 0.98
[…]
Extension to T2V Generation. To further validate the versatility of CGCE, we extend our evaluation to T2V generation. We apply our method as a plug-and-play module to several modern T2V models, including CogX-2B, CogX-5B [36] and Hunyuan [17]. For this task, we compare CGCE against the base models and two baselines: SAFREE and T2VU. As presented in Fig. 4 and Fig. 4, CGCE effectively reduces the generation of unsafe video content across all tested models. Our method consistently outperforms both SAFREE and T2VU, a method designed specifically for video concept erasure. These results demonstrate that the effectiveness of CGCE extends beyond images. Since CGCE operates on the text embedding space, it can be seamlessly applied to safeguard video generation models, highlighting its potential as a versatile, cross-modal safety solution.
4.4 Efficiency Analysis
Table 8 shows that CGCE is highly efficient in both training and inference time. The classifier training takes only 440 seconds, which is significantly faster than the roughly 4500 seconds required by fine-tuning methods like ESD and STEREO. At inference, CGCE’s runtime remains comparable to other methods. This efficiency stems from two key design choices. First, the classifier is lightweight, so its safety check adds negligible overhead to the overall generation time. Second, the iterative refinement process for unsafe prompts is fast-converging, typically requiring only a few steps. This combination of a fast one-time setup and competitive inference speed makes CGCE a practical and efficient solution for concept erasure.
Table 8: Efficiency comparison for the nudity removal task with SD-v1.4 backbone. All experiments were run on a single NVIDIA A6000 GPU, with reported times representing the average over 50 generated images using 50 denoising steps.
[…]
ESD [8] $\sim$4500 2.73 94.65
UCE [9] $\sim$1 2.73 2.23
RECE [10] $\sim$3 2.78 2.23
STEREO [30] $\sim$ 4560 2.72 94.65
SLD-Max [28] 0 2.81 0
SAFREE [38] 0 4.02 0
[…]
4.5 Ablation Study
We conduct an ablation study to analyze the effect of the step size $\eta$ used in the embedding refinement process for nudity removal task with SD-v1.4. As shown in Tab. 9, the step size $\eta$ plays a crucial role in balancing erasure effectiveness and utility preservation. As $\eta$ increases, the ASR against all adversarial prompts consistently decreases, indicating a more robust erasure. However, this comes at a cost to the model’s generative quality. For instance, increasing $\eta$ from 1.0 to 1.5 improves robustness but leads to degrade the generative quality of the model. Based on these results, we find that a step size of $\eta=1.0$ provides the best trade-off for the nudity removal task, achieving strong robustness against adversarial attacks while maintaining high utility.
Table 9: Ablation study on the step size $\eta$.
[…]
0.1 20.22 19.65 10.70 16.78 31.30
0.5 11.03 13.68 5.50 16.61 31.23
1.0 4.41 3.87 1.90 16.10 31.03
1.5 1.10 0.00 0.10 16.96 30.64
5 Conclusion
In this work, we address the challenge of concept erasure, where existing methods remain vulnerable to adversarial attacks and often degrade model utility. We introduce Classifier-Guided Concept Erasure (CGCE), a plug-and-play framework that provides robust safety for diverse T2I and T2V models without modifying their weights. CGCE uses a lightweight classifier to detect and refine unsafe text embeddings at inference time, enabling scalable removal of multiple concepts in a single process. By selectively adjusting unsafe embeddings, our method prevents harmful content generation while preserving model utility, achieving state-of-the-art robustness against red-teaming attacks across modern architectures.
Supplementary Material
6 Algorithm Details
Algorithm 1 Simultaneous Multi-Concept Embedding Refinement
1:Prompt embedding $\boldsymbol{\varepsilon}_{p}$, set of $L$ concept embeddings $\{\boldsymbol{\varepsilon}_{c_{l}}\}_{l=1}^{L}$, set of $L$ trained classifiers $\{\boldsymbol{f}_{\theta_{l}}\}_{l=1}^{L}$, step size $\eta$, probability threshold $\tau$, maximum number of iterations $K$.
2:Refined (safe) prompt embedding $\boldsymbol{\varepsilon}^{\prime}_{p}$.
[…]
4:for $k=0\to K-1$ do
5:   $\boldsymbol{G}^{(k)}\leftarrow\mathbf{0}$ $\triangleright$ Initialize aggregate gradient vector
6:   $count\leftarrow 0$ $\triangleright$ Initialize counter for detected concepts
7:   for $l=1\to L$ do $\triangleright$ Iterate over all concept classifiers
8:     $prob^{(k)}_{l},\boldsymbol{s}^{(k)}_{l}\leftarrow\boldsymbol{f}_{\theta_{l}}(\boldsymbol{\varepsilon}_{p}^{(k)},\boldsymbol{\varepsilon}_{c_{l}})$ $\triangleright$ Get probability and importance scores
9:     if $prob^{(k)}_{l}>\tau$ then $\triangleright$ Check if concept $l$ is detected
10:        $count\leftarrow count+1$
11:        $\boldsymbol{g}^{(k)}_{l}\leftarrow\boldsymbol{s}^{(k)}_{l}\odot\nabla_{\boldsymbol{\varepsilon}_{p}^{(k)}}\boldsymbol{f}_{\theta_{l}}(\boldsymbol{\varepsilon}_{p}^{(k)},\boldsymbol{\varepsilon}_{c_{l}})$ $\triangleright$ Compute weighted gradient for concept $l$
[…]
13:     end if
14:   end for
15:   if $count=0$ then
16:     break $\triangleright$ Stop refinement if no concepts are detected
17:   end if
[…]
19:end for
[…]
21:return $\boldsymbol{\varepsilon}^{\prime}_{p}$
Algorithm 1 presents the complete pseudocode for our simultaneous multi-concept embedding refinement. The algorithm iteratively refines a prompt’s text embedding $\boldsymbol{\varepsilon}_{p}$, over $K$ steps. In each step, it checks the prompt against $L$ different concept classifiers. If any classifiers detect their respective concepts, their normalized and weighted gradients are aggregated to compute a single update vector $\boldsymbol{G}^{(k)}$. This vector is then used to steer the embedding away from all detected concepts simultaneously, ensuring a safe and efficient refinement process. Notably, this algorithm inherently supports the single-concept erasure case, as it simplifies to the single-classifier update when $L=1$.
7 Additional Experimental Details
7.1 Prompt Template
To ensure training data of our classifier diversity, we utilized multiple LLMs, including Gemini 2.5 Pro [3] and Qwen2.5-7B-Instruct [31]. Each LLM was instructed with a specific system prompt for each target concept, as detailed below.
Prompt Template for Van Gogh Concept Create exactly 1000 unique pairs of prompts for image generation. Each pair must consist of: 1. A safe prompt - either with no artistic style or with a different artistic style (like Picasso, Rembrandt, Warhol, or Caravaggio). 2. A minimally modified version of the safe prompt that naturally incorporates the word “Van Gogh”. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the natural inclusion of “Van Gogh”. Examples: Safe: A man cooking in the kitchen, Picasso style. Unsafe: A man cooking in the kitchen, Van Gogh style. Safe: Students studying in library. Unsafe: Students studying in Van Gogh style library Safe: Rembrandt painting of a couple walking. Unsafe: Van Gogh painting of a couple walking.
Prompt Template for Nudity Concept Generate exactly 1000 unique prompt pairs for image generation. Each pair must consist of: 1. A safe, appropriate prompt. 2. A sexual, unsafe version that is similar but contains sexual content. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the sexual elements. Examples: Safe: A woman cooking in the kitchen. Unsafe: A nude woman cooking in the kitchen. Safe: A woman gardening in her backyard. Unsafe: A topless woman gardening in her backyard. Safe: A couple walking in the park. Unsafe: A couple having sex in the park.
Prompt Template for Church Concept Create exactly 1000 unique pairs of prompts for image generation. Each pair must consist of: 1. A safe prompt: a description that does not contain “church” anywhere. 2. A minimally modified version of the safe prompt that naturally incorporates the word “church”. The prompts within each pair should be as similar as possible in structure and content, with the main difference being the natural inclusion of “church”. Examples: Safe: A tall building with a bell tower. Unsafe: A church with a bell tower. Safe: A woman praying in a quiet place. Unsafe: A woman praying in a church. Safe: A girl playing with her dog in the park. Unsafe: A girl playing with her dog near the church in the park.
7.2 Hyperparameters
The primary hyperparameter for our refinement process is the step size $\eta$. We found that the optimal value for $\eta$ is task-specific and must be set empirically for each concept and model combination. For example, removing a broad concept like “nudity” requires a different step size than removing a specific artistic style or object on the same model. The optimal $\eta$ also varies across different model architectures for the same concept. The values used for all our experiments are detailed in Tab. 10.
Table 10: The step size $\eta$ of each model.
[…]
Van Gogh SD-v1.4 0.15
Church SD-v1.4 0.50
Nudity SD-v1.4 1.0
SD-v3 1.0
FLUX 1.5
Switti-AR 2.0
Infinity-2B 0.5
CogX-2B 2.0
CogX-5B 2.0
Hunyuan 1.0
7.3 Classifier Architecture
Our classifier is designed as a lightweight and highly efficient module that operates on the output of a pre-trained text encoder. The architecture includes an MLP for dimensionality reduction, a multi-head cross-attention layer to find concept-related tokens, and a final MLP with a sigmoid function for binary classification.
A primary design goal is a lightweight safeguard. As shown in Tab. 11, our classifiers are extremely small, with only a few million parameters, making them several orders of magnitude smaller than the text encoders they safeguard, such as CLIP-L/14 [24] (400M params), T5-XL [25] (3B params), or T5-XXL [25] (11B params). This small footprint is critical to our method’s low inference overhead.
This architecture also makes our classifiers highly reusable and model-agnostic. Because the classifier operates on the text encoder’s output, a single trained classifier can be applied to any generative model that employs the same text encoder. For example, as shown in Tab. 11, the classifier trained on T5-XXL embeddings can be used as a plug-and-play safeguard for SD-v3 [7], FLUX [18], CogX-2B, and CogX-5B [36]. This versatility extends to complex MLLMs, such as the LLaVA-Llama-3-8B [4] encoder used by Hunyuan [17], demonstrating the broad applicability of our approach.
Table 11: Classifier architecture details for each text encoder.“Embed Dim” is the text encoder’s output dimension, and “Hidden Dim” is our classifier’s internal dimension.
[…]
SD-v1.4 CLIP-L/14 768 256 0.5M
SD-v3 CLIP-L/14 + G/14 2048 512 2.4M
T5-XXL 4096 1024 9.5M
FLUX CLIP-L/14 768 256 0.5M
T5-XXL 4096 1024 9.5M
Switti-AR CLIP-L/14 + G/14 2048 512 2.4M
Infinity-2B T5-XL 2048 1024 7.4M
CogX-2B T5-XXL 4096 1024 9.5M
CogX-5B T5-XXL 4096 1024 9.5M
Hunyuan CLIP-L/14 768 256 0.5M
LLaVA-Llama-3-8B 4096 1024 9.5M
8 Multi-Concept Erasure
Figure 5: Qualitative evaluation of CGCE’s effectiveness in multi-concept erasure with FLUX. Sensitive content (*) has been masked for publication.
Figure 5 provides a qualitative evaluation of this capability on the FLUX model, using prompts that combine three concepts: an unsafe concept “nudity”, an artistic style “Van Gogh”, and an object “church”. As shown in Fig. 5, the base FLUX model generates an image containing all three concepts. When CGCE is applied with only the nudity classifier, it successfully removes the unsafe content while correctly preserving the “Van Gogh” style and the “church”. By progressively adding more classifiers, our method scales effectively. When all three classifiers are applied, CGCE successfully erases all specified concepts, demonstrating that our framework is highly scalable and can effectively remove multiple unwanted concepts simultaneously.
9 Ablation Study on Importance Score $\boldsymbol{s}$
Table 12: Ablation study on the effectiveness of using the token importance score $\boldsymbol{s}$.
[…]
$\times$ 6.98 5.39 6.99 5.26 2.20 40.14 16.67 30.99
4.62 2.47 4.41 3.87 1.90 27.46 16.10 31.03
We conduct an ablation study to validate the importance of using the token importance score $\boldsymbol{s}$ in our refinement process. As shown in Tab. 12, we compare our full method, which uses the weighted gradient ($\boldsymbol{s}\odot\nabla\boldsymbol{f}_{\theta}$), against a variant that omits the importance score (using only the raw gradient $\nabla\boldsymbol{f}_{\theta}$). The results demonstrate that the importance score is a critical component. The full method not only achieves significantly better robustness, lowering the ASR across all six benchmarks, but it also improves the model’s generative utility. This confirms that weighting the gradient by token importance is key to achieving both precise erasure and high utility preservation.
10 Extended Visual Results
Figure 6 provides further qualitative results for nudity removal on the SD-v1.4 backbone [27], comparing CGCE to other methods. Our method successfully erases the unsafe content from the generated image. Crucially, for safe, unrelated prompts, it demonstrates superior utility preservation by producing high-quality images that are identical to the original model’s output.
This finding is reinforced in Fig. 7, which evaluates the removal of the “Van Gogh” artistic style and the “church” object. CGCE effectively erases the target concepts while keeping other styles and objects unchanged. In contrast, we observe that while a robust method like STEREO [30] also removes the target concept, it suffers from significant utility degradation. For example, when the STEREO model (which was fine-tuned to erase “church”) is prompted for an unrelated object such as a “guitar” or “telephone”, it incorrectly generates an image of a “car”.
We also demonstrate the generalization of CGCE by applying it to different T2I backbones. Figure 8 and Figure 9 provide visualizations of nudity removal on modern DiT-based models (SD-v3 [7], FLUX [18]) and VAR-based models (Switti-AR [33], Infinity-2B [11]). CGCE successfully removes the unsafe content while preserving the surrounding context and other safe objects in the image.
Finally, our method extends seamlessly to T2V generation. We provide qualitative results for nudity removal on CogX-2B (Fig. 10, Fig. 11), CogX-5B [36] (Fig. 12, Fig. 13), and Hunyuan [17] (Fig. 14, Fig. 15). CGCE effectively removes the unsafe content across video frames while maintaining the temporal consistency and visual integrity of other objects and actions. This contrasts with other video-erasure methods like T2VU [37], which can generate noticeable visual artifacts.
Figure 6: Qualitative evaluation of CGCE and other concept erasure methods on the SD-v1.4 backbone. The figure compares performance on nudity erasure, evaluating each method’s ability to remove the target concept while preserving unrelated ones. Sensitive content (*) has been masked for publication.
Figure 7: Qualitative evaluation of CGCE and other concept erasure methods on the SD-v1.4 backbone. The figure compares performance on “Van Gogh” style erasure (left) and “Church” object erasure (right), evaluating each method’s ability to remove the target concept while preserving unrelated ones.
Figure 8: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with SD-v3 (left) and FLUX (right). Sensitive content (*) has been masked for publication.
Figure 9: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with Switti-AR (left) and Infinity-2B (right). Sensitive content (*) has been masked for publication.
Figure 10: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with CogVideoX-2B. Sensitive content (*) has been masked for publication.
Figure 11: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with CogVideoX-2B. Sensitive content (*) has been masked for publication.
Figure 12: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with CogVideoX-5B. Sensitive content (*) has been masked for publication.
Figure 13: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with CogVideoX-5B. Sensitive content (*) has been masked for publication.
Figure 14: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with HunyuanVideo. Sensitive content (*) has been masked for publication.
Figure 15: Qualitative evaluation of CGCE’s effectiveness in erasing nudity concepts, compared to baseline methods with HunyuanVideo. Sensitive content (*) has been masked for publication.